# Examining Input Modalities and Visual Feedback Designs in Mobile Expressive Writing

### v1

@@ Abstract.
Expressive writing is an established approach for stress management, and recent practices include information technology. Although mobile interfaces have the potential to support daily stress management practices, interface designs for such mobile expressive writing and their effects on stress relief still lack empirical understanding. To fill the gap, we examined the interface design of mobile expressive writing by investigating the influence of input modalities and visual feedback designs on usability and perceived cathartic effects through in-the-wild studies. While our studies confirmed the stress relief effects of mobile expressive writing, our results offer important insights in interface design. We found keyboard-based text entry more user-friendly and preferred over voice messages due to its privacy friendliness and reflection process. Participants expressed different reasons for preferring different post-writing visual feedback depending on the cause and type of stress. This paper also discusses future research opportunities in interface designs for mobile expressive writing.
@@ 1. Introduction
Excessive mental stress has a negative impact on many aspects of life, including mental and physical well-being (Veit and Ware, 1983; Hammen, 2005; Mushtaq et al., 2018). Due to its high global prevalence, several approaches to lower stress have been developed, primarily in the field of psychology. One established method is expressive writing, referring to the disclosure of distressing experiences and their related emotions typically by journaling (Pennebaker, 1997; Pennebaker and Chung, 2007). Expressive writing can provide people with a sense of stress relief (Pennebaker and Chung, 2007), known as a cathartic effect (Breuer and Freud, 1955). Expressive writing started with pen and paper (Pennebaker and Chung, 2007), but the practice has expanded to the information technology (Hardey, 2002). More recently, mobile apps (Wang et al., 2018; Wu and Bailey, 2018; MacIsaac et al., 2022) and chatbots (Howe et al., 2022; Park et al., 2021) have also been investigated to support expressive writing. For example, smartphone apps or chatbots ask users to type their negative emotions (Howe et al., 2022; Abd-Alrazaq et al., 2021; Ahmed et al., 2022). Smartphones, in particular, are now highly ubiquitous, which creates an opportunity to investigate and distribute affordable, effective, and easily accessible stress-lowering tools to those suffering from stress.
Due to recent technological advances, speech input is now a viable input option on smartphones. Previous HCI work has found that speech input has an advantage in speed (Ruan et al., 2018), workload (Rzepka et al., 2022), and content richness (Luo, 2022) over keyboard input. Further, studies about expressive writing without smartphones or chatbots have found that verbal expression had a more positive effect than written expression (e.g. (Esterling et al., 1994; Murray and Segal, 1994)). Speech-enabled expressive writing with smartphones is currently underexplored, as most existing expressive writing support systems rely on keyboard text entry (see e.g. (Howe et al., 2022; Park et al., 2021)). This warrants further studies into the suitability of speech input in expressive writing.
Another powerful aspect of smartphones is their high programmability, allowing developers to build experiences that mimic those outside the digital world. Clinical psychology has shown that deleting objects tied to emotion helps to resolve negative moods (Briñol et al., 2013; Lee and Schwarz, 2011; Soesilo et al., 2021). People may write a note with their negative emotion and put it into a shredder or simply tear it up by hand (Nussbaum, 1995; Soesilo et al., 2021). This consequent action is a widely recognized practice to obtain an explicit feeling of removing or destroying the negative experience (Briñol et al., 2013; Lee and Schwarz, 2011). Similar features can be implemented as digital mobile experiences through different visual feedback solutions, but the potential of these solutions has not well examined.
This paper aims to better understand the cathartic effect of expressive writing on mobile devices. We first examined the impact of speech input versus typed input on user experience and the cathartic effect within a mobile expressive writing application. We also adapted key findings from the first investigation and improved the mobile solution. We then compared different visual feedback designs that are based on prior work and were provided immediately after the users engaged in expressive writing. Specifically, these studies answer two research questions:
@@
How do different visual feedback designs in mobile expressive writing influence stress-relieving effect, and the perceived usability and preference?
This work conducts two in-the-wild studies (15 and 23 days long with 28 and 36 participants, respectively) using a mobile chatbot application and offers the following contributions:
Investigation of the effects of input modalities and visual feedback designs provided after emotional disclosure in mobile expressive writing.
Design and research implications based on our two studies, highlighting future research directions in technology-supported expressive writing and stress coping.
@@ 2. Related Work
Our work is contextualized within the intersection of digital tools for expressive writing (Pennebaker, 1997; Pennebaker and Chung, 2007) and related HCI aspects.
2.1. Expressive Writing
The therapeutic use of writing is called expressive writing (Pennebaker, 1997; Pennebaker and Chung, 2007) (also called written emotional disclosure (Frisina et al., 2004) or therapeutic writing (Wright and Chung, 2001)). In expressive writing, people write freely about their traumatic or stressful events with their emotions in detail. Undergoing this expressive writing session is known for its various benefits in the long term, ranging from physical and mental health to social and behavioral aspects (Pennebaker, 1997; Reinhold et al., 2018). Expressive writing can bring increased distress and negative mood right after writing as short-term effects (Esterling et al., 1999) while prior research also found a decrease in depressive symptoms (Reinhold et al., 2018). To unfold these complex and contradictory outcomes, there has been a large discussion on the process underlying expressive writing.
The mechanism of how expressive writing works has not been fully understood, but there are three common theories to explain: inhibition theory, exposure theory, and cognitive processing (Reinhold et al., 2018). Early work on expressive writing followed inhibition theory, assuming that inhibiting thoughts or emotions caused a high mental workload, and therefore, expressing thoughts and emotions led to physical and mental health (Pennebaker and Beall, 1986). Another explanation, exposure theory, proposed that expressive writing might work as an imaginative exposure to negative emotional events and let people inured to stimuli. During exposure, the response to the stress was unlearned or replaced by other responses, and this process might relate to the outcome of expressive writing (Sloan and Marx, 2004). Cognitive processing spotlighted the content of expressive writing and claimed that the process of constructing a coherent story helped people organize their experiences and thoughts and change perspectives (Pennebaker and Chung, 2007). Considering these theories, expressive writing lets us confront and unleash our negative experiences and emotions, which we tend to inhibit or avoid. Although it may temporarily provoke negative moods, it still helps people restructure their cognition toward stressful events and brings various benefits.
@@ 2.1. Expressive Writing
Along with psychology revealing its effects and mechanisms, existing research has explored the applications of information technology to support expressive writing (Sauter, 2014; Hardey, 2002). Experiments with EmotionDiary showed how the tool was most effective for people with past experience on emotional disclosure on social networking services (Lee et al., 2016). Another group designed chatbot interactions for expressive writing, and the responsive chatbot enhanced emotional disclosure than writing on a document (Park et al., 2021). Other applications support journaling (Wang et al., 2018), reflection (McDuff et al., 2012) or recollection (Peesapati et al., 2010) of everyday emotional experiences as well as trauma and stress experiences.
One previous study explored chatbot designs to mitigate workplace stress by comparing three types of interventions: taking a break, supporting mindfulness, and emotional expression (Howe et al., 2022). While the emotional expression intervention reduced stress most efficiently, participants rated it the lowest because of two drawbacks: it required the highest effort, and they felt more stressed while reflecting their stress and the associated negative emotion. Technology-supported expressive writing, like conventional longhand writing, helped disclose emotions and reduce stress, while short-term deterioration of emotions was observed as well, and its burden became problematic. Therefore, there is a need to explore usability aspects of interface designs for mobile expressive writing for cathartic effect.
2.2. The Effect of Input Modalities in Expressive Writing
Previous work found that verbal expressions (e.g., speaking into a tape recorder) could improve cognitive change and enhance immune function more (Esterling et al., 1994) than written expressions (e.g., writing on a piece of paper) while both verbal and written expressions led to increase in negative emotion (Murray and Segal, 1994). For the degree of emotional expression, there are two conflicting results; one reported greater expression in written mode (Esterling et al., 1994) while the other in a verbal condition (Murray and Segal, 1994). Prior work has also examined the effects of verbal and written forms only in self-disclosure but also in disclosure to others. Iacovelli et al. (Iacovelli and Johnson, 2012) found that face-to-face disclosure showed stronger physiological improvements than remote communication via text messaging. Other work has compared therapy interviews (in verbal form with therapists) with expressive writing (in written form by oneself) and found that therapy interviews helped a shift to positive affect (Donnelly and Murray, 1991; Murray et al., 1989) and cognitive change (Murray et al., 1989). Combining these findings, verbal expression, which is the most natural and prevailing modality for therapy, may have a greater effect on post-expression mood and stress reduction than written expression.
Many of the existing expressive writing technologies and studies have mainly employed text entry using keyboards (either physical or virtual) (Abd-Alrazaq et al., 2019; Park et al., 2021; Howe et al., 2022). Brewin and Lennard’s study on comparing typing against writing for a long term revealed that typing resulted in less negative effects but less self-rated disclosure (Brewin and Lennard, 1999). This result motivates more extensive explorations on the effect of input modalities in technology-enabled expressive writing. Speech-based input is in particular a promising text entry approach commonly available in various smartphones and mobile devices (Enge, 2020). Speech-based input has been shown as superior over keyboard-based methods for entering short messages on touchscreen mobile devices (Ruan et al., 2018). It also has demonstrated benefits in information search with conversational agents (Rzepka et al., 2022) and food journaling (Luo, 2022). However, its effect on mobile expresstive writing is still underexplored. Comparative studies in input modalities thus contribute to deepening the understanding of the effect of technology-enabled expressive writing.
@@ 2.2. The Effect of Input Modalities in Expressive Writing
2.3. Feedback Effects in Expressive Writing
Prior work in psychology has shown that additional actions on the expressed negative emotion, such as physically throwing it away, can enhance the sense of stress relief (Briñol et al., 2013; Lee and Schwarz, 2011; Soesilo et al., 2021). These studies show that the physical clearing of materialized emotions leads to the psychological clearing of emotions. Washing or sanitizing hands can also diminish the extent and impact of previous negative sensations (Xu et al., 2012) or feelings of guilt (Zhong and Liljenquist, 2006). These psychological effects depend on the intensity of the physical action or the degree of embodiment. Soesilo et al’s experiment found that shredding the paper with scissors reduced envy more than tearing by hand (Soesilo et al., 2021). Briñol et al. (Briñol et al., 2013) also tested the effect of interaction on a computer, and revealed that moving files with negative emotions to the trash with a mouse had a greater effect than merely imagining them. As our cognition is based on sensory organs, the mental state is grounded in the physical state (Barsalou, 2008; Lee and Schwarz, 2021). Therefore, the embodiment and agency of the elimination action are important, but the effects when others perform the elimination movements are not fully clarified (Lee and Schwarz, 2021).
Digital technology offers great flexibility on the consequent actions on expresstive writing. Grieger et al. (Grieger et al., 2021) demonstrated that in virtual reality, punching objects displaying stressful messages promoted relaxation and encouraged self-reflection. Björling et al.(Björling et al., 2019) developed a virtual reality system to reduce stress in teenagers, which allowed users to burn their negative thoughts with leaves in a fire. While previous studies in virtual reality explored the effects of eliminating emotional objects, most studies and applications on mobile devices have implemented no feedback or only text feedback (Park et al., 2021; Howe et al., 2022). Yet, the effects of the visual feedback after emotional disclosure on mobile devices remain relatively underexplored.
@@ 3. Study 1: Examination of Input Modalities
We first conducted a study to explore the influence of using different input methods to answer RQ1. The following study protocol was approved by the Institutional Review Board of the first author’s institution.
@@ 3.1. System Implementation
We developed a chatbot system that allows users to perform expressive writing on their smartphones (Figure 1). We adopted chatbot interaction as it can enhance emotional disclosure rather than writing on a document (Park et al., 2021). We chose LINE as the underlying platform because it is one of the most widely used messaging apps in Japan, the country where we conducted our studies. In our implementation, participants could express their stressful experience that happened that day and how they felt then, using a given input method (Figure 1b). The instructions were based on the standard wording (Pennebaker and Chung, 2007), but translated into more casual, conversational words in reference to the study that designed a chatbot for expressive writing (Park et al., 2021). After the system received stress-related emotional expression content, it asked participants to rate the perceived stress changes at that moment (Figure 1e). The system could also send a message to remind participants and encourage them to share their experiences. We prepared fixed messages for notifications and responses. We prepared fixed notification and response messages for the chatbot, which also deliberately included emojis as prior work had shown the effect of making a chatbot more credible (Beattie et al., 2020) or enjoyable (Fadhil et al., 2018).
[Two screenshots of the Chatbot system interface we used in Study 1.]Two screenshots of the Chatbot system interface we used in Study 1. The first screenshot with areas labeled (a) to (c), and another screenshot with areas labeled (d) and (e). Section (a) dispalys the general chat UI of the system, showing the message sent from the system on the left side and message users sent on the right side. Section (b) displays the message sent by the system saying: ”Good evening. How was your day? Please tell us about a stressful event that happened today, and how you felt about it. No matter how trivial the event may seem, we ask you to describe that event using text input only. Please avoid using voice-note or other input methods. You may input your response by clicking on the keyboard symbol in the lower-left corner of the screen. We kindly ask that you submit your response by 24:00. Thank you for your cooperation.” Section (c) displays a menu at the bottom of the screen with two buttons: ”Evaluate your Stress Level” with a five-star icon, and ”Describe a Stressful Event” with a pen and paper icon. Below these buttons is text saying, ”* These buttons are only available if you have responded to our most recent survey.” and a clickable link saying, ”Click here to input your response”, followed by a ”Optional Menu” button. Section (d) shows a reply message to the expressive writing that says, ”Thank you for sharing this stressful event. We understand that some of this information may be difficult to share, but we greatly appreciate your cooperation. We kindly ask you to evaluate your current stress level once more.” Section (e) shows a message for ”Stress Evaluation” with buttons on it. The message starts with ”Comparing your responses before and after you described a stressful event, how do you feel your stress level has changed? Please press the button that corresponds with your response only one time.” Below this text are five buttons labeled ”Greatly Increased” in red, ”Increased” in light red, ”No Change” in gray, ”Decreased” in light green, and ”Greatly Decreased” in green. A note at the bottom reads, ”* As only responses received by 24:00 will be considered valid, we ask for your cooperation in responding as soon as possible. * Please make your selection only one time. We ask for your patience as our system may take some time to respond to you.”

### v3

@@ Abstract.
Expressive writing is an established approach for stress management. Recently, information technologies, such as smartphones, have also been explored for expressive writing. Although mobile interfaces have the potential to support various daily writing activities, interface designs for mobile expressive writing and their effects on stress relief still lack empirical understanding. We examined the interface design of mobile expressive writing by investigating the influence of input modalities and visual feedback designs on usability and perceived cathartic effects through field studies. While our studies confirmed the stress-relieving effects of mobile expressive writing, our results offer important insights into interface design. We found keyboard-based text entry more suited and preferred over voice messages for its privacy and reflective nature. Participants expressed different reasons for preferring different post-writing visual feedback depending on the cause and type of stress. This work advances interface design for mobile expressive writing and deepens understanding of its effects.
@@ 1. Introduction
Expressive writing is an established approach for mental well-being management, referring to textual disclosure of stressful experiences and their related emotions (Pennebaker, 1997; Pennebaker and Chung, 2007). It can provide people with a sense of stress relief (Pennebaker and Chung, 2007), known as a cathartic effect (Breuer and Freud, 1955). Expressive writing originated from pen and paper (Pennebaker and Chung, 2007). But the practice has expanded to information technologies (Hardey, 2002), including mobile apps (Wang et al., 2018; Wu and Bailey, 2018; MacIsaac et al., 2022) and chatbots (Howe et al., 2022; Park et al., 2021) as mental well-being is becoming an important issue to the general user populations. Digital implementations maintain many of the traditional benefits of expressive writing while adding advantages of social sharing, such as engagement, reflection, and guidance (Park et al., 2021). While digital implementations of expressive writing have demonstrated cathartic effects (Howe et al., 2022), there still exists a research gap to fully leverage the capability of recent information technology. In particular, as smartphones are now highly ubiquitous, optimizing their use can offer an opportunity for people to perform expressive writing at hand, promoting routine stress management and contributing to their mental well-being.
Despite its promising potential, there is a lack of understanding in mobile interface designs for expressive writing. Regarding the input modalities, existing research has revealed advantages of speech-based input for mobile interfaces in speed (Ruan et al., 2018), workload (Rzepka et al., 2022), and content richness (Luo, 2022). However, it remains underexplored how such an input modality could contribute to user experience in mobile expressive writing. Prior work also shows that verbal expression may provoke a sense of releasing stress and enhance cognitive change (Esterling et al., 1994; Murray et al., 1989), but it is unknown how speech-based expressive writing compares to traditional text-based input with respect to cathartic effects and user experience. Feedback to users after disclosure is another key design component for mobile expressive writing. Clinical psychology has shown that deleting objects tied to emotion helps to resolve negative moods (Briñol et al., 2013; Lee and Schwarz, 2011; Soesilo et al., 2021). People may write a note with their negative emotion and put it into a shredder or simply tear it up by hand (Nussbaum, 1995; Soesilo et al., 2021), which can provide an explicit feeling of removing or destroying the negative experience (Briñol et al., 2013; Lee and Schwarz, 2011). Translating this mechanism into mobile interfaces presents a design challenge: balancing the emotional impact with the level of user effort required. To address this, we introduced the concept of passive intervention (Zhao et al., 2023) and compared it with interactive visual feedback and a condition offering only textual responses. Our exploration can inform future design strategies for technology-supported emotional regulation.
This paper aims to better understand the cathartic effect of expressive writing on mobile devices. More specifically, we design our studies to answer the following two questions:
@@
How do different visual feedback designs in mobile expressive writing influence the stress-relieving effect, and the perceived usability and preference?
To answer these questions, we conducted two field studies (15 and 23 days long with 28 and 36 participants, respectively). Our first study examined the effect of keyboard-based and speech-based input, and the second experiment focused on the effect of post-disclosure interaction. Our studies highlighted the unique benefits of each of the input modalities and post-disclosure interaction design instead of the superiority of one of those. Our results suggest future research directions in mobile expressive writing. The contributions of this paper, therefore, are two-folded:
Empirical evidence on how input modalities and post-disclosure visual feedback designs influence the cathartic effect and user experience in mobile expressive writing, along with a discussion on underlying psychological processes.
Design and research implications based on our two studies, highlighting the necessity of considering not only stress levels but also stress types in the design of technology-supported expressive writing and stress coping.
@@ 2. Related Work
2.1. Expressive Writing: Outcomes, Mechanisms, and Protocols
Writing can have a therapeutic effect under certain conditions. The therapeutic use of writing, known as expressive writing (Pennebaker, 1997; Pennebaker and Chung, 2007), is also referred to as written emotional disclosure (Frisina et al., 2004) or therapeutic writing (Wright and Chung, 2001). Pennebaker et al. introduced expressive writing as a therapeutic method for PTSD and found it improved physical health (Pennebaker and Beall, 1986). Since then, research has revealed various benefits, including reduced anxiety (Reynolds et al., 2000) and depression (Soliday et al., 2004), cognitive and self-esteem changes (Pennebaker, 1997; Murray and Segal, 1994), decreased aggression (Kliewer et al., 2011), and improved physical conditions (Pennebaker and Beall, 1986). However, while expressive writing promotes well-being, it may also initially increase distress and negative mood (Esterling et al., 1999), though studies suggest a subsequent reduction in depressive symptoms (Reinhold et al., 2018). These complex and sometimes contradictory outcomes have led to extensive discussions on the underlying processes of expressive writing.
Three common theories attempt to explain the effects of expressive writing: inhibition theory, exposure theory, and cognitive processing (Reinhold et al., 2018). Early research followed inhibition theory, suggesting that suppressing thoughts and emotions increases mental workload, and therefore, expressing them improves physical and mental health (Pennebaker and Beall, 1986). Exposure theory posits that expressive writing functions as imaginative exposure to negative emotional events, helping individuals become desensitized. Through this process, people unlearn past stress responses and adopt new ones, potentially explaining expressive writing’s effects (Sloan and Marx, 2004). Cognitive processing focuses on content, arguing that constructing a coherent narrative helps individuals organize experiences, shift perspectives, and gain clarity (Pennebaker and Chung, 2007; Niles et al., 2016). These theories suggest that expressive writing encourages confronting and releasing suppressed emotions. While it may temporarily intensify negative moods, it ultimately helps restructure cognition around stressful events, yielding various benefits.
Researchers have also explored the boundary conditions of expressive writing to deepen understanding. In Pennebaker’s basic paradigm, individuals write about their deepest thoughts and feelings regarding traumatic or stressful experiences (Pennebaker and Chung, 2007). Participants are assured confidentiality and encouraged to focus on emotions rather than spelling, structure, or grammar. Each session lasts 15–30 minutes, repeated over 3–5 days. Building on this protocol, research has examined various influencing factors (Pennebaker and Chung, 2007; Smyth and Pennebaker, 2008), expanding the scope of expressive writing. Notably, it is not limited to traumatic or negative experiences but can include positive events (Burton and King, 2004; King and Miner, 2000) or others’ traumatic episodes (Greenberg et al., 1996). Session duration also varies; even two-minute writing sessions over two days have shown health benefits (Burton and King, 2008). Additionally, session frequency is flexible, with a single 10–15 minute session with positive framing or a one-hour intensive session yielding similar benefits to the traditional multi-day approach (Chung and Pennebaker, 2008; Smyth and Pennebaker, 2008).
2.2. Technology-Supported Expressive Writing
@@
Research has explored the use of information technology to support expressive writing (Hardey, 2002; Sauter, 2014). Experiments with EmotionDiary, an expressive writing application for online social networking services (SNS), revealed that it was particularly effective for individuals with prior experience in emotional disclosure on SNS (Lee et al., 2016). Another study designed chatbot interactions for expressive writing, showing that a responsive chatbot enhanced emotional disclosure more than writing on a document (Park et al., 2021). Other applications support journaling (Wang et al., 2018), reflection (McDuff et al., 2012), and recollection (Peesapati et al., 2010) of both everyday emotional experiences and trauma-related stress. These technologies can maintain the benefits associated with traditional expressive writing while providing potential advantages typically afforded by social sharing, such as enhanced engagement, reflection, and guidance, without incurring the risks often associated with human-based social disclosure, including stigma and loss of privacy (Park et al., 2021).
Research also compared different chatbot interventions for workplace stress: taking breaks, promoting mindfulness, and emotional expression (Howe et al., 2022). Though emotional expression was most effective in reducing stress, it was rated lowest due to the high effort required and increased stress during reflection. Technology-supported expressive writing, revealed emotions and reduced stress, but caused short-term emotional decline and was burdensome. Thus, there is still a need to improve mobile expressive writing interfaces for cathartic benefits.
2.2.1. The Effect of Input Modalities in Expressive Writing
Verbal forms of expressive writing, so-called expressive talking (Harrist et al., 2007; Langer et al., 2012), yield unique outcomes. Previous research indicates verbal expressions (e.g., speaking into a tape recorder) may enhance cognitive change and immune function more effectively than written expressions (e.g., writing on paper) (Esterling et al., 1994), or at least offer comparable health benefits (Murray and Segal, 1994; Lyubomirsky et al., 2006). The expression mode also influences word choice: Talking involves more words (Murray and Segal, 1994), fewer negative emotional words (Esterling et al., 1994), and a more coherent narrative (Newton, 2015) than writing. In expressive writing, overuse of negative emotional words leads to negative results, while creating a coherent story offers advantages (Pennebaker and Chung, 2007); these linguistic features may be linked to unique outcomes.
Many expressive writing technologies use physical or virtual keyboards (Abd-Alrazaq et al., 2019; Park et al., 2021; Howe et al., 2022). Brewin and Lennard found that typing, compared to pen and paper, led to fewer negative effects but also less self-disclosure (Brewin and Lennard, 1999). This highlights the need to explore the impact of different input methods. Speech-based input is promising on smartphones (Enge, 2020), outperforming keyboards for short messages (Ruan et al., 2018), aiding search with conversational agents (Rzepka et al., 2022), and assisting in food journaling (Luo, 2022). Research on verbal disclosure versus written self-disclosure suggests essential differences. For example, Iacovelli et al. observed stronger physiological benefits from face-to-face disclosure compared to text messaging (Iacovelli and Johnson, 2012). Similarly, comparisons between therapy interviews (with therapists) and expressive writing (self-administered) indicate that interviews promote positive affect (Donnelly and Murray, 1991; Murray et al., 1989) and cognitive change (Murray et al., 1989). These projects indicate the potential for further exploration of alternative input modalities, particularly speech, within the context of mobile expressive writing.
@@ 2.2.1. The Effect of Input Modalities in Expressive Writing
2.2.2. Feedback Effects in Expressive Writing and Materializing Emotional Catharsis Method
Researchers have studied feedback’s role in enhancing expressive writing, which is inherently self-directed (Pennebaker and Beall, 1986; Pennebaker and Chung, 2007). Feedback has been introduced post-writing, especially for learners with disabilities (Gersten and Baker, 2001; MacArthur et al., 1996), using input from teachers (Sawyer et al., 1992) or peers (Wong et al., 1996; O’Leary et al., 2018). AI-driven conversational feedback, like Park et al.’s Responsive Chat, encourages self-awareness and reflection by posing questions on emotions and impacts (Park et al., 2021). However, user privacy is a major issue in adaptive feedback (Pennebaker and Chung, 2007). While technology that offers a low-cost, privacy-preserving feedback solution exists (Bond and Pennebaker, 2012; Park et al., 2021), feedback designs that do not engage in the content of disclosure would liberate developers and users from privacy concerns.
Psychological research highlights the cathartic effect of materializing emotions as tangible entities and symbolically destroying them. Prior work has shown that additional actions on expressed negative emotions, such as physically discarding them, can enhance stress relief (Briñol et al., 2013; Lee and Schwarz, 2011; Soesilo et al., 2021; Kanaya and Kawai, 2024). Briñol et al.(Briñol et al., 2013) found that even digital interactions, such as dragging files representing negative emotions to the trash, had a greater effect than merely imagining their disposal. Digital technology further extends this concept by enabling flexible expressive writing and interactive emotional processing. In virtual reality, physically punching objects with stressful messages promoted relaxation(Grieger et al., 2021), while Björling et al.(Björling et al., 2019) developed a VR system where teenagers burned negative thoughts as leaves to reduce stress. However, most mobile device studies focus on limited feedback(Park et al., 2021; Howe et al., 2022), and the impact of visual feedback after emotional sharing remains underexplored.
There has been ongoing debate regarding the key factors that contribute to the effectiveness of materialized emotional disposal. Lee et al. (Lee and Schwarz, 2021) reported that physical engagement enhanced cathartic effects. According to grounded cognition theory (Barsalou, 2008), intense physical movements reinforce the perceived meaning of the action, facilitating a deeper sense of de-fusion (i.e., decouple) from negative emotions (Hatvany et al., 2018). From this perspective, interactive interventions that promote a sense of agency (Moore, 2016) on disposal may be especially effective in mobile interface design. In contrast, Briñol et al. (Briñol et al., 2013) found that the interpretation of disposal may play a more critical role than the sensorimotor experience itself. According to this view, simply presenting the image of emotional disposal could be sufficient, even when the system carries out the action rather than the user. However, the psychological effect of such system-initiated disposal remains underexplored (Lee and Schwarz, 2011). Additionally, prior research suggests that passive interventions, which do not require active user engagement, can be effective in high-stress contexts (Zhao et al., 2023). Given that minimal-effort interventions are preferable in everyday stress management applications (Howe et al., 2022), it is vital to investigate how system-driven, passive interventions differ in psychological impact from interactive ones.
@@ 3. Study 1: Examination of Input Modalities
As discussed in the previous section, while voice input may facilitate emotional disclosure and cognitive processing, prior studies have mainly focused on keyboard input, leaving its effects unexplored. To address this gap and answer RQ1, we conducted a user study investigating how different input modalities affect the cathartic effect and usability in expressive writing. This section introduces the design of our first user study. The following study protocol was approved by the Institutional Review Board of the first author’s institution.
@@ 3.1. System Implementation
We developed a smartphone-based system for expressive writing (Figure 1). We chose text chat interaction over document writing to enhance emotional disclosure (Park et al., 2021). LINE was selected as the platform due to its popularity in Japan, where our study was conducted. Participants shared daily stress experiences and feelings using a specific input method (Figure 1b). After completing their writing, they rated perceived stress changes, indicating the momentary cathartic effect (Figure 1e). The system also sent reminders. We predefined all messages, using emojis to increase credibility (Beattie et al., 2020) and enjoyment (Fadhil et al., 2018).